### v1

# Just Because You Can, Doesn’t Mean You Should: LLMs for Data Fitting

Abstract
Large Language Models (LLMs) are being applied in a wide array of settings, well beyond the typical language-oriented use cases. In particular, LLMs are increasingly used as a plug-and-play method for fitting data and generating predictions. Prior work has shown that LLMs, via in-context learning or supervised fine-tuning, can perform competitively with many tabular supervised learning techniques in terms of predictive performance. However, we identify a critical vulnerability of using LLMs for data fitting – making changes to data representation that are completely irrelevant to the underlying learning task can drastically alter LLMs’ predictions on the same data. For example, simply changing variable names can sway the size of prediction error by as much as 82% in certain settings. Such prediction sensitivity with respect to task-irrelevant variations manifests under both in-context learning and supervised fine-tuning, for both close-weight and open-weight general-purpose LLMs. Moreover, by examining the attention scores of an open-weight LLM, we discover a non-uniform attention pattern: training examples and variable names/values which happen to occupy certain positions in the prompt receive more attention when output tokens are generated, even though different positions are expected to receive roughly the same attention. This partially explains the sensitivity in the presence of task-irrelevant variations. We also consider a state-of-the-art tabular foundation model (TabPFN) trained specifically for data fitting. Despite being explicitly designed to achieve prediction robustness, TabPFN is still not immune to task-irrelevant variations. Overall, despite LLMs’ impressive predictive capabilities, currently they lack even the basic level of robustness to be used as a principled data-fitting tool.
1 Introduction
Large language models (LLMs), as one of the most representative Generative AI applications, have been adopted across a wide variety of domains and tasks. While the majority of LLMs’ use cases are expectedly language-related (translation, Q&A, writing, ideation, etc.), there are also increasing uses of LLMs for tasks not apparently associated with language. We focus on one such practice, where LLM is used as an “off-the-shelf” data-fitting tool to generate predictions of interest. Data fitting refers to the task of using sets $\{(\boldsymbol{X},Y)\}$ of numerical input values $\boldsymbol{X}$ and target values $Y$ to learn (fit) a model $f(.)$ – i.e., $Y\sim f(\boldsymbol{X})$ – that can generate accurate predictions given inputs. For example, LLMs have been used for predicting property prices based on property features (e.g., number of bedrooms, bathrooms, etc.) and location information Tanlamai et al., (2024); performing time series forecasting and numeric prediction tasks in general by treating input data as a sequence of string tokens Gruver et al., (2023); Jin et al., (2023); Das et al., (2024); Goswami et al., (2024); Vacareanu et al., (2024); detecting anomalies in time series data Alnegheimish et al., (2024); Dong et al., (2024); Liu et al., (2025); generating synthetic data in data-hungry domains Seedat et al., (2023); Wang et al., (2024); Brand et al., (2023); Isomura et al., (2024); and support causal effect estimation by predicting counterfactual outcomes (Huynh et al.,, 2025). In all of these examples, LLMs essentially act as a drop-in replacement for traditional, tabular supervised learning techniques to generate predictions based on input data (See SI Appendix Table S1 for more details about these examples).
A priori, it is unclear whether LLMs should be used for data fitting and predictions. On one hand, hailed as a general-purpose technology Eloundou et al., (2024), there is evidence that LLMs’ impressive capabilities extend beyond language tasks and can perform data fitting quite well. Prior work has shown that LLMs, with just a few input-output examples included in the prompt (i.e., using few-shot prompting), can outperform many tabular supervised learning techniques for regression tasks (Vacareanu et al.,, 2024). Furthermore, when feature names are semantically meaningful, LLMs can perform feature selection automatically, which may further benefit data fitting (Jeong et al.,, 2024). On the other hand, LLMs vs. tabular supervised learning techniques perform “data fitting” and “prediction” in fundamentally different ways. Tabular supervised learning techniques perform data fitting by following well-defined algorithmic procedures that fit parametric functions (e.g., linear regression) or non-parametric functions (e.g., random forest) to training data. The trained models then produce predictions by directly applying the learned function (i.e., plug in the feature values of new input data). In contrast, the vast majority of LLMs today are based on the transformer architecture, which auto-regressively generates the next token based on prior tokens. With this architecture, LLMs perform data fitting by (i) having training examples in the prompt, namely in-context learning Brown et al., (2020), or by (ii) supervised fine-tuning using the training data Hu et al., (2021). [^1] We do not consider the use case where LLMs generate and/or execute code to build tabular models. In such case, LLMs are being used as a coding tool rather than a data-fitting tool. They generate predictions one token at a time, where each token may correspond to only a segment of the prediction value. As an illustrative example, consider how an LLM generates a numerical prediction value. To generate a prediction value of 0.1875, OpenAI’s GPT-4o model needs to generate 4 tokens precisely in the following order: “0”, “.”, “187”, and “5”. [^2] This example can be replicated at: https://platform.openai.com/tokenizer These fundamental differences raise the question whether LLMs provide a principled or robust way to fit numerical data and obtain predictions.
In this work, we offer a critical assessment of using LLMs for data fitting. We construct a numeric prediction task and create synthetic data from scratch. Instead of using publicly available datasets, which LLMs may have been exposed to during training, using synthetic data enables a clean evaluation of LLM’s data-fitting capabilities. We experiment with general-purpose LLMs that are both close-weight (GPT-4o-mini) and open-weight (Llama-3-8B-instruct) as well as special-purpose tabular foundation models (TabPFN, Hollmann et al.,, 2025). Consistent with prior work Gruver et al., (2023); Vacareanu et al., (2024), we find that LLMs perform reasonably well (especially after supervised fine-tuning), achieving performance that is competitive with commonly-used tabular supervised learning techniques.
However, just because LLMs can be used for data fitting does not mean they should be. A principled data-fitting technique not only needs to achieve reasonable predictive performance, but also needs to possess at least a basic level of robustness so as not to be affected by factors that are fundamentally unrelated to the task itself. Our analyses show that LLMs have alarmingly poor prediction robustness. In particular, we apply several different types of task-irrelevant variations – such as altering variable names, variable (column) order, example (row) order, and data format – none of which has any effect on the underlying data-fitting task. While canonical tabular supervised learning techniques are completely unaffected (essentially by their design), LLMs’ predictions for the same data can change dramatically, causing the size of prediction error to sway as much as 82% in some settings. Further statistical tests confirm that such prediction sensitivity is indeed attributable to the task-irrelevant variations and cannot be explained by randomness in the LLM token generation process.
We also investigate the attention scores in an open-weight LLM to explore the potential sources of LLMs’ poor prediction robustness. By summing up the attention scores from output tokens (which make up the predicted value) to each of the in-context training examples, we discover a phenomenon of “U-shaped” attention distribution, where the first and last training examples receive more attention than examples in the middle. Moreover, this phenomenon also manifests within in-context examples – variable names and values located towards the beginning or end of a given example receive more attention than other positions. Such uneven attention distributions allow training examples and variables that happen to occupy the beginning / end positions in a prompt to have unduly large influence, thereby contributing to substantial variations in LLMs’ predictions.
Besides general-purpose LLMs, there are growing efforts to build specialized foundation models for tabular data (“tabular LLM” or “tabular foundation models”, Sui et al.,, 2024; Hollmann et al.,, 2025). We repeat our evaluations on a state-of-the-art tabular foundation model, TabPFN Hollmann et al., (2025), which adopts a modified attention mechanism as an attempt to facilitate the invariance of predictions with respect to row order and variable order. Nonetheless, we find that it is still not immune to task-irrelevant variations.
Our results reveal a fundamental challenge of using LLMs for data fitting: prediction robustness with respect to task-irrelevant variations, which is completely taken for granted when using tabular supervised learning techniques, can hardly be guaranteed in LLMs; at best, it can only be partially engineered in special-purpose foundation models. As an analogy, using LLMs for data fitting currently is like using a calculator for adding numbers, where the calculator produces different answers when given the same set of numbers in a different order. Therefore, we caution against the increasingly prevalent use of LLMs as a black-box data-fitting tool – the prediction quality cannot be ensured due to lack of robustness with respect to task-irrelevant variations.
Beyond data fitting, it is worth highlighting that our findings also have broader implications for current LLMs as a paradigm of learning. We take the data-fitting task not as the ultimate objective per se but as an opportunity to better understand the limitations of LLMs. The ability to disentangle what is relevant vs. irrelevant for a task is widely recognized as a fundamental requirement for forming abstraction (Giere,, 2010; Weisberg,, 2012), learning conceptual categories (Rosch,, 2024), cognitive processing (e.g., the selective attention theory Broadbent,, 1958; Johnston and Dark,, 1986), all of which are important in problem-solving. When LLMs change their predictions in response to task-irrelevant variations, it calls into question whether LLMs possess basic “competence” for the data-fitting task. Together with recent work showing that inconsequential changes in prompting can significantly affect LLMs’ performance on solving math problems (Mirzadeh et al.,, 2024) and imitating human behaviors (Gao et al.,, 2025), we highlight that the use of current LLMs as principled problem-solving tools requires careful reconsideration.
2 Materials and Methods
Because modern LLMs have been pre-trained with massive datasets collected from the Internet, it is inappropriate to evaluate LLMs’ data-fitting capabilities on publicly available datasets which may have already been exposed to the LLMs during training. Therefore, we create synthetic data from scratch for all experiments in this paper. Specifically, we simulate the following linear data generation process for $p$ input variables:
$Y=\sum_{i=1}^{p}\beta_{i}X_{i}+\varepsilon$ (1)
In the baseline setting, $p=10$ and each input variable $X_{i}$ is randomly and independently sampled from either $N(0,1)$ or $\exp(1)$; coefficients $\{\beta_{1},\ldots,\beta_{p}\}$ are sampled from $Unif[0,1]$ and fixed; and $\varepsilon$ is randomly sampled from $N(0,0.1^{2})$. We choose such a simple linear data-generation process for two reasons. First, its parametric form gives us fine-grained control over different aspects of the data-fitting task, such as number of input variables ($p$) and distributional characteristics of these variables. Second, it represents an “easy” data-fitting task. If LLMs exhibit problematic data-fitting capabilities under this setup, it is more likely due to a fundamental limitation of current LLMs rather than difficulty of the data-fitting tasks. We use this data-generation process to simulate a training dataset of 4,000 instances and a testing dataset of 1,000 instances.
We focus on two specific data-fitting capabilities. We first evaluate the performance level, i.e., how well LLMs fit data and generate predictions compared to tabular supervised learning techniques. This determines whether LLMs have the capability to fit data accurately. More importantly, we also evaluate prediction sensitivity, i.e., whether LLM-generated predictions are robust with respect to changes in the inputs that are irrelevant to the data-fitting tasks (and are extraneous for tabular supervised learning techniques). This sheds light on whether LLMs should be used for data fitting.
We carry out comprehensive data-fitting experiments on GPT-4o-mini (a representative general-purpose close-weight LLM), [^3] Among many close-weight commercial LLMs, GPT-4o-mini is one of the most cost-effective flagship LLMs that support both completion and fine-tuning. Llama-3-8B-instruct (a representative general-purpose open-weight LLM), as well as TabPFN (currently best-performing tabular foundation model specifically trained for data-fitting tasks, Hollmann et al.,, 2025). We perform LLM-based data fitting via both in-context learning (ICL) and supervised fine-tuning (SFT). For GPT-4o-mini, we use OpenAI’s batch processing API for ICL and fine-tuning API for SFT. For Llama-3-8B-instruct, we download the model from Hugging Face [^4] Source: https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. and carry out ICL using resources provided by Minnesota Supercomputing Institute. For the tabular foundation model (TabPFN), we follow TabPFN documentation to perform ICL using its Python implementation. The detailed prompts and procedures for both ICL and SFT are reported in SI Appendix, “LLM Prompts and Prediction Procedures” Section. Given the objective nature of the task, we use greedy decoding for all LLMs (i.e., selecting the next token as the one receiving highest predicted probability). This is done by setting LLM generation temperature to 0 and disabling random token sampling whenever possible. We also fix the random seed during SFT of LLMs as well as the training of all tabular supervised learning models to ensure reproducibility and fair comparisons across different models. All models are evaluated on the same 1,000 testing instances.
3 LLMs Can Achieve Competitive Prediction Accuracy
3.1 Performance Level
Starting with GPT-4o-mini, we evaluate the LLMs’ predictive performance level as compared to six widely-used tabular supervised learning techniques, including linear regression (the oracle model in this case), LASSO regression, $k$-NN, Support Vector Regression (SVR), Random Forest, and Multi-Layer Perception (MLP). Each supervised technique is trained on the same training data of 4,000 instances, with hyper-parameters tuned via 5-fold cross-validation. For reference, we also include a naive benchmark, namely using the mean of outcome variable on the training data as (constant) prediction. For GPT-4o-mini under SFT, we use all 4,000 training instances for fine-tuning. Under ICL, following the common practice in few-shot prompting, we include either 10 or 20 instances sampled from training data into the task prompt (denoted as ICL10 and ICL20). We also explore an ICL prompt with 500 instances (denoted as ICL500), representing a large-scale setting that is still within the context window of GPT-4o-mini. We evaluate predictive performance on the testing data using the standard Mean Absolute Error (MAE) metric. The results are shown in Table 1.
Table 1: MAE Comparisons between LLM and Supervised Learning Techniques (GPT-4o-mini)
$X_{i}\sim N(0,1)$ $X_{i}\sim\exp(1)$
Naive Benchmark 1.5331 1.5869
Linear Regression (Oracle) 0.0792 0.0791
LASSO 0.0825 0.0827
SVR 0.0792 0.0791
Random Forest 0.5574 0.4848
$k$-NN 0.5414 0.5416
MLP 0.0944 0.0905
ICL10 0.9246 0.8749
ICL20 0.9283 0.7717
ICL500 0.9315 0.7830
SFT 0.1219 0.1313
Not surprisingly, linear regression and other tabular supervised learning techniques that can sufficiently approximate a linear relationship (LASSO, SVR, and MLP) all perform quite well on this task, whereas Random Forest and $k$-NN perform comparatively worse. More importantly, LLM-based data fitting achieves respectable performance. With SFT, LLM outperforms Random Forest and $k$-NN and catches up with the performance of the other techniques. With ICL, even with only 10 or 20 examples (substantially smaller “training” size than the other models), it still meaningfully outperforms the naive benchmark, indicating that LLM can conduct few-shot learning to some degree in a prediction task. Further increasing the number of examples in ICL to 500 does not necessarily improve predictive performance. [^5] We further consider a few variations to the above baseline setting, including different numbers of variables, variable scales, and the presence of outliers in data. Overall, we find that LLMs achieve reasonable performance levels across these variations, especially with SFT.
4 LLMs Have Poor Prediction Robustness with Respect to Task-Irrelevant Variations
Achieving advantageous predictive performance is not the only requirement for data-fitting techniques in practice. We now evaluate LLMs’ prediction sensitivity – whether (and how much) LLMs’ predictions for the same data instances would vary in the presence of task-irrelevant variations to input data, i.e., when we apply changes that do not alter the underlying prediction task in any meaningful way. More specifically, we consider five types of task-irrelevant variations, described as follows:
  • Variable Names: change the names of input variables from “X0”, “X1”, $\ldots$, “X9” to “First_Variable”, “Second_Variable”, $\ldots$, “Tenth_Variable”;
  • Variable Order: randomly shuffle the order of variables (but maintain the same variable order for each data instance);
  • Number of Digits after decimal point: round each numerical value from 15 digits (the default precision of a double float number in Python) to 10 digits. [^6] While doing so technically reduces the precision of input variables, the applied changes are highly negligible as compared to how we measure and report predictive performance (4 digits after decimal point). We therefore still consider it as a task-irrelevant change.
  • Format: change the natural language format of data instance representation (e.g., “A data point has X1 0.1269…”) to a JSON key-value format.
  • Row Order: for ICL, reverse the order of the few-shot examples in LLM prompts; for SFT, use a different random seed (which would lead to a different ordering of data instances during fine-tuning).
Note that these changes do not meaningfully alter the actual data or the data-fitting task, and tabular supervised learning techniques would not be materially affected by them. However, we observe that LLM-generated predictions have poor robustness across all of these changes, and the sensitivity can be quite high under some settings. In particular, let $MAE_{base}$ and $MAE_{change}$ respectively denote the predictive performance before and after applying task-irrelevant changes, we define the prediction sensitivity metric as
$\text{\%Sensitivity}=\frac{|MAE_{base}-MAE_{change}|}{MAE_{base}}\times 100\%$ (2)
and we report the results in Table 2.
Table 2: LLM Prediction Sensitivity with Respect to Task-Irrelevant Variations (GPT-4o-mini)
$X_{i}\sim N(0,1)$
ICL10 ICL20 ICL500 SFT
Variable Names 26.06% 28.00% 29.47% 2.19%
Variable Order 3.46% 0.08% 11.68% 0.07%
Number of Digits 1.64% 5.74% 0.26% 0.07%
Format 5.70% 7.86% 5.78% 23.78%
Row Order 34.00% 25.87% 9.77% 5.12%
$X_{i}\sim\exp(1)$
ICL10 ICL20 ICL500 SFT
Variable Names 82.06% 39.68% 5.76% 2.39%
Variable Order 1.37% 8.64% 10.53% 3.18%
Number of Digits 0.33% 8.52% 11.95% 5.19%
Format 3.40% 0.43% 4.27% 3.59%
Row Order 69.19% 32.05% 1.95% 8.34%
All task-irrelevant variations lead to nontrivial sensitivity in LLM predictions, which can be as high as 82% under ICL and 24% under SFT. Worse yet, we do not observe any clear patterns across different settings – SFT does not necessarily lead to more robust predictions than ICL; ICL with more in-context examples does not lead to greater robustness; and sensitivity also varies with the distribution of input variables. In other words, LLM-based data fitting and predictions not only suffer from poor robustness with respect to fundamentally task-irrelevant variations, the degree of sensitivity is also hard to anticipate a priori.
In addition to applying each task-irrelevant change separately, we also consider the compounding impact of more than one task-irrelevant change. As an illustrative example, we examine a combination of “Variable Order” and “Format”, where the training data are represented in a JSON key-value format and each instance has randomized variable order. This poses no effect on tabular supervised learning techniques, but we again observe clear prediction sensitivity in LLM. With exponentially distributed covariates, GPT-4o-mini performance varies by 8.44% under ICL10 setting, 1.48% under ICL20 setting, 6.51% under ICL500 setting, and 0.33% under SFT setting. Furthermore, sensitivity due to the combined change differs from that of the individual changes in an unpredictable manner (bigger in some settings and smaller in others).
4.1 Prediction Sensitivity Cannot Be Attributed to Randomness
Even though we set LLM generation temperature to 0 in all of the experiments, re-running the exact same prompt may still generate slightly different predictions, due to ties in predicted logit values of candidate tokens. [^7] For instance, if two candidate tokens (e.g., “123” and “124”) share the same predicted logit value during one inference step, some LLMs would randomly choose one of these two candidates as the output token. OpenAI API does not offer precise control over this randomness. We perform statistical tests to ascertain that the prediction sensitivity we observe is not attributed to such randomness in token generation.
Specifically, to assess the “baseline” prediction sensitivity that is purely caused by token generation randomness, we re-run each setting (both ICL and SFT) for a second time, without applying any task-irrelevant changes. For a given testing instance $\boldsymbol{x}$, we would obtain two predictions from the two identical runs, $LLM(\boldsymbol{x})$ and $LLM^{\prime}(\boldsymbol{x})$, that could potentially be slightly different. For the same testing instance under a given task-irrelevant change, let $LLM^{change}(\boldsymbol{x})$ denote the LLM’s prediction. Next, we compute prediction variation due to randomness $\Delta_{rand}=|LLM(\boldsymbol{x})-LLM^{\prime}(\boldsymbol{x})|$ as well as prediction variation due to the task-irrelevant change $\Delta_{change}=|LLM(\boldsymbol{x})-LLM^{change}(\boldsymbol{x})|$.
If a task-irrelevant change leads to prediction sensitivity that cannot be simply explained away by token generation randomness, we would expect $\Delta_{change}$ to significantly exceed $\Delta_{rand}$ on average. We check this by performing one-sided paired $t$-tests comparing the two quantities across 1,000 testing instances. We find that the differences are statistically significant at 0.001 level for all but two settings (where the prediction sensitivity is small); the only exceptions are SFT with normally distributed variables under variable name change ($p=0.036$, significant at 0.05 level) and under number of digits change ($p=0.064$, significant at 0.1 level).
4.2 Prediction Sensitivity Cannot Be Avoided by Prompting
We further demonstrate that prediction sensitivity cannot simply be prevented by telling the LLM not to be sensitive with respect to task-irrelevant variations. Let’s consider the Row Order variation as an illustrative example. Under the ICL setting (where training examples are part of the prompt), we insert an extra instruction after the ICL examples that specifies “The presented $K$ examples above are not in any particular order.” This is intended to signal to the LLM that the specific ordering of training examples has no bearing on the data-fitting task.
In Table 3, we present the predictive performance and sensitivity associated with three configurations: (i) Baseline (i.e., the original ICL prompting); (ii) Row Order (i.e., applying the row order change); (iii) Row Order ++ Extra Instruction. We report results based on exponentially distributed variables (the results based on normally distributed variables are qualitatively consistent).
Table 3: Predictive Performance and Sensitivity with Extra Prompt Instruction ($X_{i}\sim\exp(1)$, prediction sensitivity compared to Baseline reported in parentheses)
ICL10 ICL20 ICL500
(i) Baseline MAE 0.8749 0.7717 0.7830
(ii) Row Order 1.4803 1.0190 0.7983
(69.19%) (32.05%) (1.95%)
(iii) Row Order ++ Extra Inst. 1.2602 0.9498 0.8532
(44.03%) (23.09%) (8.96%)
The extra prompt instruction is clearly ineffective in countering prediction sensitivity induced by row order change. In configuration (iii), we still observe nontrivial sensitivity (up to 44.03%). Moreover, comparing configurations (ii) and (iii), we can see that inserting the extra instruction, which is intended to improve prediction robustness, can actually lead to higher prediction sensitivity in some settings (e.g., ICL500).
In general, prompt engineering alone cannot resolve LLMs’ prediction sensitivity with respect to task-irrelevant variations. Even if one can somehow anticipate the variations to guard against ahead of time and insert corresponding prompt instructions, doing so can end up increasing (rather than decreasing) prediction sensitivity.
5 Replication with an Open-Weight LLM and Mechanism Exploration
We repeat the performance level and sensitivity analyses with a flagship open-weight model, Llama-3-8B-instruct. The open-weight nature of this LLM also provides a unique opportunity to explore the underlying reason for LLMs’ lack of prediction robustness, by examining the internal attention scores. We focus on the ICL10 and ICL20 settings here, because ICL500 or SFT with this model require extensive computing resources that are beyond what is available to us. We report results under exponentially distributed variables (the results under normally distributed variables are qualitatively similar).
5.1 Results
Table 4 summarizes both the predictive performance and the prediction sensitivity of Llama-3-8B-instruct. Overall, it achieves even better predictive performance than GPT-4o-mini under the ICL settings. However, like GPT-4o-mini, its predictions are sensitive with respect to task-irrelevant variations. Note that we set the Llama model to use greedy decoding without random token sampling, so re-running the same setting yields exactly the same set of predictions, thus the observed prediction sensitivity still cannot be attributed to token generation randomness.
Table 4: Performance Level and Sensitivity of Llama-3-8B-instruct ($X_{i}\sim\exp(1)$)
ICL10 ICL20
Baseline MAE 0.7911 0.6757
Variable Names 2.82% 0.80%
Variable Order 5.28% 9.75%
Number of Digits 6.38% 5.49%
Format 0.64% 3.32%
Row Order 9.68% 3.67%
5.2 Attention-Based Mechanism Exploration
The black-box nature of LLMs makes it challenging to pin down exactly why they generate some specific tokens over others. Several research studies have explored the use of attention scores to derive possible explanations for certain LLM behaviors (e.g., Elhage et al.,, 2021; Zhou et al.,, 2024; Kamath et al.,, 2025). For auto-regressive LLMs, attention scores are crucial in determining which token to generate next Vaswani et al., (2017). We investigate the cross-attention from the tokens that make up the prediction value to the tokens that represent the ICL examples, in order to probe the potential mechanism behind LLMs’ prediction sensitivity.
More concretely, consider an ICL prompt with $K$ examples included in it. Let $P_{k}$ denote the set of tokens that make up the $k$-th example, let $O$ denote the output tokens that make up the LLM’s prediction value. For each pair of tokens $o\in O$ and $p\in P_{k}$, let $Attn(o,p)$ denote the cross-attention score that token $o$ pays to token pp. [^8] Llama-3-8B-instruct has 32 attention heads. $Attn(o,p)$ is the average attention score across all attention heads. We can define a “row attention” that quantifies the amount of attention from output tokens to one ICL example:
$RowAttn(k)=\frac{1}{|O|}\sum_{o\in O}\sum_{p\in P_{k}}Attn(o,p)$ (3)
We normalize it by token count in the output so that row attention is not skewed by output length. Moreover, $P_{k}$ can be broken down to three subsets of tokens – those representing variable names, those representing variable values, and other miscellaneous tokens (blank space, comma, etc.). Let $PName_{ik}$ and $PVal_{ik}$ respectively denote the variable name tokens and variable value tokens corresponding to the $i$-th variable ($i\in\{1,\ldots,10\}$) in the $k$-th example. We further define “variable name attention” and “variable value attention” as follows:
[…]
In a canonical data-fitting task – learning a model that maps input values to target predictions – there should not a priori be any differential focus on specific rows / variable names / variable values (e.g., variable names are not even part of the canonical definition of a data-fitting task). In other words, the expectation would be that $RowAttn(k)$ is roughly the same across $k$; $NameAttn(i)$ and $ValAttn(i)$ should be roughly the same across $i$. In fact, these types of expected attention “regularities” are explicitly engineered into special-purpose tabular foundation models (e.g., TabPFN, Hollmann et al.,, 2025, as will be discussed later). However, as can be seen in the results in Figure 1, we observe a “U-shaped” attention pattern: rows at the beginning or end of the ICL examples receive more attention than other rows, and variable names / values at the beginning or end of a given example receive more attention than those at other positions. Such attention distributions are statistically significant – a Kruskal-Wallis test shows that all three attention distributions are different from being uniform ($p<0.001$ in all three cases).
Figure 1: Attention Distributions by Rows, Variable Names, and Variable Values (Llama-3-8B-instruct under ICL10 setting)
We observe similar “U-shaped” attention distributions under the ICL20 setting (with 20 ICL examples that are different from those in ICL10). This further indicates that the “U-shaped” pattern is associated with specific positions in the ICL prompt rather than the training examples we use. See SI Appendix, “Additional Analyses of Attention Distributions” Section for the detailed results.
The uneven attention distributions represent a plausible explanation for LLMs’ prediction sensitivity, because training examples and variables that happen to occupy the beginning / end positions in a prompt end up having larger influence in determining the output tokens. To illustrate this, we repeat the ICL10 experiment but shuffle the variable order. The “U-shaped” attention distributions persist, and variables that show up at beginning / end of each example by random chance now receive greater attentions from the output tokens. We include the results in SI Appendix, “Additional Analyses of Attention Distributions” Section.
Our findings are reminiscent of the phenomena of “position bias” (Zhao et al.,, 2021; Cobbina and Zhou,, 2025, where positions of in-context examples in user prompts can influence few-shot learning performance), “attention sink” (Xiao et al.,, 2023, where attention scores toward initial tokens tend to be stronger even if those tokens are not semantically meaningful), and “lost in the middle” (Liu et al.,, 2023, where LLMs’ information retrieval performance tends to be worse when relevant information is in the middle part of a long context). Meanwhile, the attention pattern has a richer, more nuanced structure in our context, as the “U-shaped” distributions manifest both across few-shot examples and within each example (across variable names / values).
6 Specialized Tabular Foundation Models Are Not Immune to Prediction Sensitivity
Next, we extend our analyses to tabular foundation models specifically designed for data-fitting and prediction tasks. TabPFN Hollmann et al., (2025) represents a state-of-the-art tabular foundation model that has outperformed supervised learning techniques on a wide variety of benchmarking datasets. Importantly, TabPFN recognizes that data fitting should be invariant with respect to variable order and row order, and explicitly tries to facilitate this by implementing a “two-way attention” mechanism where each tabular cell attends to other features in the same row as well as to other values under the same column.
TabPFN can carry out ICL with up to 500 input variables and 10,000 data instances. Therefore, in addition to the ICL10 / ICL20 / ICL500 settings, we also put the entire training dataset (4,000 training instances, denoted as ICL4000) to TabPFN and obtain predictions on the same 1,000 testing instances. Note that TabPFN does not take variable names and requires a standard dataframe format, so we focus on the other three task-irrelevant variations, namely “Variable Order”, “Number of Digits”, and “Row Order”. The computational efficiency of TabPFN allows us to repeat the evaluations 100 times, each with different training / testing datasets simulated from the same data-generation process. We report the mean and standard deviation of performance level and prediction sensitivity in Table 5 for exponentially distributed variables (normally distributed variables yield similar results).
Table 5: Performance Level and Sensitivity of TabPFN ($X_{i}\sim\exp(1)$, standard errors of MAE and prediction sensitivity are reported in parentheses)
ICL10 ICL20 ICL500 ICL4000
Baseline MAE 1.0917 0.7282 0.1048 0.0974
(0.2712) (0.1905) (0.0098) (0.0081)
Variable Order 2.87% 3.34% 2.54% 1.94%
(0.0368) (0.0310) (0.0196) (0.0150)
Number of Digits 1.14% 1.60% 0.86% 0.75%
(0.0095) (0.0141) (0.0061) (0.0062)
Row Order 0.74% 1.51% 0.73% 0.67%
(0.0099) (0.0130) (0.0063) (0.0058)
TabPNF achieves impressive predictive performance – with all 4,000 training examples, its average MAE clearly outperforms general-purpose LLMs (GPT-4o-mini and Llama-3-8B-instruct) under both ICL and SFT. However, it is still not immune to task-irrelevant variations, with average prediction sensitivity ranging from 0.67% to 3.34%. Like the Llama model, re-running TabPFN with the same prompt yields exactly the same set of predictions, so the observed prediction sensitivity again cannot be attributed to token generation randomness.
Although the degree of prediction sensitivity may seem small, we stress that task-irrelevant variations should not cause any prediction changes to begin with. As mentioned earlier, this could be viewed as essentially analogous to having a calculator that produces different results when adding up the same set of numbers in a different order. The fact that TabPFN is not immune to task-irrelevant variations points to a deeper issue. Despite TabPFN’s explicit architectural choices to “build in” the invariance with respect to variable order and row order, it cannot fully achieve the said invariance. In contrast, supervised learning techniques are completely invariant to these changes by design. We argue that this highlights a fundamental challenge with using current LLMs for data fitting: although robustness with respect to some task-irrelevant variations could be facilitated to a limited extent with special-purpose tabular foundation models, it generally cannot be guaranteed.
7 Conclusions
Modern LLMs have demonstrated impressive performance across a wide array of settings. As a result, they are increasingly being used as “general-purpose” AI tools, beyond the typical language-oriented tasks. One such task is data fitting and prediction, where LLMs have been shown to perform competitively as or even outperform common tabular supervised learning techniques via in-context learning or supervised fine-tuning (Gruver et al.,, 2023; Vacareanu et al.,, 2024; Alnegheimish et al.,, 2024; Dong et al.,, 2024).
We argue that just because LLMs can be used for data fitting does not mean they should be. Through extensive numerical experiments, we demonstrate that current LLMs have poor prediction robustness with respect to task-irrelevant variations. Altering variable names, order of variables, order of training examples, or inconsequential details of data representation – all of which are completely irrelevant to the underlying data-fitting task – can significantly change the LLMs’ predictions on the same data.
Such prediction sensitivity persists across different learning methods (in-context learning or supervised fine-tuning) as well as different LLMs (general-purpose close- or open-weight LLMs and special-purpose tabular foundation models), and cannot simply be attributed to randomness in token generation. Instead, we provide descriptive evidence that the output tokens are paying more attention to training examples and variable names / values that happen to occupy the beginning or end positions of a prompt, even if those positions are not more important a priori. The non-uniform attention distributions provide a plausible explanation behind spurious sensitivity in LLMs’ predictions.
We believe the absence of even a basic level of prediction robustness brings several challenges. First, it introduces additional complications to the (already difficult) challenge of LLM interpretability (Nanda et al.,, 2023; Conmy et al.,, 2023). Why an LLM produces a certain prediction depends not only on the underlying data but also on specific, task-irrelevant word choices in the prompt. Second, it raises safety, reliability, and trust concerns when LLM-based data fitting and prediction systems are deployed in practice. For instance, if predictions on the same data can be easily altered by changing some variable names, how much would we want to trust those predictions, let alone to make (potentially high-stakes) decisions based on them? Last but not least, people familiar with tabular supervised learning techniques typically take for granted that task-irrelevant variations would not affect their models or predictions. The idiosyncratic choices of LLM users related to data representation can greatly change the results of the same analyses. As one of the consequences, it can further worsen the reproducibility crisis in science (Baker,, 2016; Youyou et al.,, 2023) due to the ever increasing demand for data analysis, data fitting, and predictive modeling in scientific inquiry and the increasing use of LLMs for this purpose.
Our study resonates with recent work that reveals LLM limitations in imitating human behaviors (Gao et al.,, 2025, where the language used in prompting can significantly alter LLM behavior patterns) and mathematical reasoning (Mirzadeh et al.,, 2024, where changes to person names appearing in a math question clearly affects LLM performance). In all of these studies, the tasks of interest are commonly understood to involve learning proper abstractions and applying principled procedures, which is incompatible with observations that innocuous, task-irrelevant modifications end up causing drastic changes in LLM outputs. Taken together, this line of inquiry highlights an important challenge of current LLMs, namely their inability to distinguish what is completely irrelevant in a given task. As discussed earlier, we believe such limitations cast serious doubt on whether current LLMs, as a learning paradigm, exhibit “competence” as a problem-solving tool. Addressing such limitations represents an important direction for future work.
Appendix A Details about Recent Work that Uses LLMs for Data-Fitting
Table 6 summarizes recent work that uses LLMs for data fitting.
Citation Data-Fitting Task How LLMs Perform Data-Fitting and Prediction
Tanlamai et al., (2024) Predict property prices based on property features (e.g., number of bedrooms / bathrooms) and location information (e.g., zip code). Zero-shot prediction. The LLM is given the persona of a realtor, and property features are given in a JSON key-value format in user prompt. The LLM is then asked to estimate the property price based on its “knowledge of similar markets and home values in the area”.
Gruver et al., (2023); Jin et al., (2023); Das et al., (2024); Goswami et al., (2024) Time series forecasting. Time series data are treated as strings of digits. Zero-shot prediction (Gruver et al.,, 2023), few-shot in-context learning (Jin et al.,, 2023), or training special-purpose time series foundation models from scratch (Das et al.,, 2024; Goswami et al.,, 2024).
Vacareanu et al., (2024) Linear and non-linear regressions. In-context learning with 500 examples.
Alnegheimish et al., (2024); Dong et al., (2024); Liu et al., (2025) Anomaly detection from time series data, either by predicting the indices of anomalous observations or by conducting time series forecasting to identify deviations from observed values as anomalies. Zero-shot prediction (Alnegheimish et al.,, 2024) and in-context learning (Dong et al.,, 2024; Liu et al.,, 2025).
Brand et al., (2023); Seedat et al., (2023); Wang et al., (2024); Isomura et al., (2024) Data augmentation, e.g., imputing missing values based on complete data records or generating new (synthetic) data based on observed data. Zero-shot prediction (Brand et al.,, 2023), In-context learning (Seedat et al.,, 2023; Wang et al.,, 2024) and supervised fine-tuning (Isomura et al.,, 2024).
Huynh et al., (2025) Causal inference, specifically by predicting counterfactual outcomes to support causal effect estimation. In-context learning.
Table 6: Details about Recent Work that Uses LLMs for Data-Fitting
Appendix B LLM Prompts and Prediction Procedures
We provide details about LLM prompts and prediction procedures used in our data-fitting experiments. For both in-context learning (ICL) and supervised fine-tuning (SFT) approaches, we use 10 variables ($p=10$) as the base configuration.
For ICL with GPT-4o-mini, we employ a structured prompt consisting of two main components: a task description and few-shot examples. The task description instructs the model: “Your job is to predict the target value based on some features. You will be given 10 features in total, including X0, X1, X2, X3, X4, X5, X6, X7, X8, X9. Please output the target value as a number. It is very important to only output the target number and nothing else.” This is followed by few-shot examples presented in the format: “You will be given a total of $K$ examples. Here is example $i$: - A data point has X0 [VAL], X1 [VAL], X2 [VAL], X3 [VAL], X4 [VAL], X5 [VAL], X6 [VAL], X7 [VAL], X8 [VAL], X9 [VAL]. The correct target value of this data point is [VAL].” This pattern is repeated for K1K-1 additional examples, with [VAL] placeholders filled with actual numerical values from the training data. During ICL prediction, each of the 1,000 test instances is presented without their ground-truth target values as user queries to obtain model completions.
For ICL with Llama-3-8B-instruct, we utilize the model’s native chat template format with specific system and user message delimiters. The prompt follows the same structure of task description and few-shot examples, supplemented with Llama-3’s special tokens to properly delineate the user-assistant conversation flow. The complete prompt is as follows:
<|start_header_id|>system<|end_header_id|> Your job is to predict the target value based on some features. You will be given 10 features in total, including X0, X1, X2, X3, X4, X5, X6, X7, X8, X9. Please output the target value as a number. It is very important to only output the target number and nothing else. <|eot_id|> <|start_header_id|>user<|end_header_id|>Predict the target for: X0 [VAL], X1 [VAL], X2 [VAL], X3 [VAL], X4 [VAL], X5 [VAL], X6 [VAL], X7 [VAL], X8 [VAL], X9 [VAL], <|eot_id|> <|start_header_id|>assistant<|end_header_id|> [VAL] <|eot_id|> (repeat for the other few-shot examples) <|start_header_id|>user<|end_header_id|>Predict the target for: X0 [VAL], X1 [VAL], X2 [VAL], X3 [VAL], X4 [VAL], X5 [VAL], X6 [VAL], X7 [VAL], X8 [VAL], X9 [VAL], <|eot_id|> <|start_header_id|>assistant<|end_header_id|>
The inference is conducted using computational resources from the Minnesota Supercomputing Institute, with each of the 1,000 test instances presented without their ground-truth target values as user queries to obtain model completions.
For SFT with GPT-4o-mini, we construct fine-tuning data points by combing the task description with individual training instances. A total of 4,000 such fine-tuning data points are provided to OpenAI’s SFT service to create a specialized fine-tuned model. The resulting fine-tuned model employs the same task description as its system instruction. During prediction with the SFT approach, each of the 1,000 test instances is presented to the fine-tuned model as a user query to generate predictions.
Appendix C Additional Analyses of Attention Distributions
Figure 3 presents the attention distributions associated with the ICL20 setting. Figure 3 presents the attention distributions associated with the ICL10 setting with randomly shuffled variable orders. The variable order after the random shuffling is ‘X4’, ‘X0’, ‘X2’, ‘X8’, ‘X5’, ‘X6’, ‘X9’, ‘X1’, ‘X3’, ‘X7’.
Figure 2: Attention Distributions by Rows, Variable Names, and Variable Values (Llama-3-8B-instruct under ICL20 setting)
Figure 3: Attention Distributions by Rows, Variable Names, and Variable Values (Llama-3-8B-instruct under ICL10 setting with randomly shuffled variable order)